OpenStudy (anonymous):

What is the slope and y interception of the line 4x+7y=21 (Need to show work)

jigglypuff314 (jigglypuff314):

Hello @erik101 and Welcome to OpenStudy! :) it would be easier to identify the slope and y-intercept if we had the equation in the slope intercept form y = mx + b so let's first try to get y alone on one side subtract 4x from both sides then divide both sides by 7 can you do that? :)

OpenStudy (anonymous):

Okay I got y=-4x+3

jigglypuff314 (jigglypuff314):

close! :D y = (-4/7) x + 3 then we have y = m x + b as the formula where m = slope and b = y-intercept

OpenStudy (anonymous):

slope -4/7x and y-intercept= 3?

jigglypuff314 (jigglypuff314):

slope is the number without that x so slope = -4/7 but yes, that's correct! :)